# Minerva: A File-Based Ransomware Detector

### v2

@@ Abstract.
This paper presents Minerva, a novel robust approach to ransomware detection. Minerva is engineered to be robust by design against evasion attacks, with architectural and feature selection choices informed by their resilience to adversarial manipulation. We conduct a comprehensive analysis of Minerva across a diverse spectrum of ransomware types, encompassing unseen ransomware as well as variants designed specifically to evade Minerva. Our evaluation showcases the ability of Minerva to accurately identify ransomware, generalize to unseen threats, and withstand evasion attacks. Furthermore, Minerva achieves remarkably low detection times, enabling the adoption of data loss prevention techniques with near-zero overhead.
@@
ccs: Security and privacy Malware and its mitigationccs: Computing methodologies Machine learning
@@ 3. Threat Model
Trusted Components. In our threat model we assume Minerva’s components to be trusted, meaning that the ransomware processes cannot tamper with their functionality. In particular, it is assumed that the Disk Activity Monitor module, and the File-Based Behavioral Detector module, can work correctly and independently of any ransomware infecting the system. These assumptions are in line with prior work in the field (Continella et al., 2016; Mehnaz et al., 2018; Kharraz and Kirda, 2017; Elkhail et al., 2023).
@@ 4.1. Minerva Architecture
From a high-level view, Minerva is comprised of two separate modules: (1) Disk Activity Monitor, and (2) File-Based Behavioral Detector. In what follows, we explain how these two modules work.
@@
The Disk Activity Monitor module is a shim that provides system-wide, real-time information on all file system operations performed. We implemented the module as a driver that uses the I/O Request Packet data structure provided by Windows systems to transparently log filesystem operations system-wide. Each filesystem-related system call performed by a process can result in many lower-level operations. The Disk Activity Monitor captures and filters all these low-level operations before processing them into feature vectors, which are then forwarded to the File-based Behavioral Detector module of Minerva. To minimize overhead, we logically divide files into critical and non-critical: critical files encompass all user files that are high-priority targets for ransomware, while non-critical files comprise system and configuration files that can easily be recovered after an attack. The Disk Activity Monitor excludes any operation performed on non-critical files and computes per-file feature vectors for all other opened files.
@@
Figure 1 illustrates the architecture of the File-based Behavioral Detector module of Minerva. This module comprises an ensemble of machine-learning classifiers trained to identify ransomware activity by analyzing the behavioral profile of files. For each critical file in the system, Minerva computes several features based on the operations received by the file within a predefined time window. These features define that file’s behavioral profile for that time window, which the Detector uses to determine whether the file has undergone malicious activity within the window. To withstand variations in ransomware behavior, such as adaptive attacks aimed at slowing ransomware activity to span multiple windows, Minerva employs a multi-tier architecture. Each tier implements a distinct fully-connected Deep Neural Network classifier trained on feature vectors computed over progressively longer time windows: low-tier classifiers utilize feature vectors computed over short time windows (e.g., $0.25$ sec), while high-tier classifiers utilize feature vectors calculated over longer windows (e.g., $4$ sec). At the end of each time window, Minerva queries the classifier of the corresponding tier to detect malicious behavioral profiles. In order to confirm malicious activity on a file, Minerva requires that at least one of its classifiers reports $K$ consecutive windows with detected malicious activity. If a classifier reports malicious activity for $K$ consecutive windows, all processes that were active on the specific file during those windows are flagged as ransomware processes. This procedure is carried out for each active critical file in the system that is receiving read or write operations.
@@ 4.2. Feature Analysis
Minerva leverages seven distinctive file-level features that accurately describe the behavioral profile of files. We carefully selected these features following an analysis of a dataset containing both benign and ransomware processes to validate our initial intuition behind the choice. Our experimental evaluation demonstrates the significant challenge ransomware faces in modifying its behavior to successfully evade all of these features, given the contrastive nature between different features.
@@
The number of read and write operations that a file undergoes within a given time window offers essential context to weigh the significance of the other features. Benign processes commonly execute occasional file operations of relatively small size over their lifetime, leading to file-based behavioral profiles characterized by a low number of operations within each time window. This is apparent from Figures 2(b) and 2(f), which respectively depict the distribution and CDF of the number of operations received by a file over 1-second intervals. Certain traditional ransomware favor using a limited number of large read and write operations during the encryption process, likely aiming to maximize I/O performance. Conversely, the majority of traditional ransomware utilize numerous small operations on the file, presumably to minimize the likelihood of detection. Similarly, in an effort to closely imitate the behavior of benign processes, evasive multiprocess ransomware imposes restrictions on the number of operations performed by each individual ransomware process and on the size of each operation (De Gaspari et al., 2022a; Zhou et al., 2023). The restriction imposed on individual operation size forces evasive multiprocess ransomware to use a greater overall number of read/write operations, distributing them across the various ransomware processes. In terms of file-based behavioral profiles, this behavior leads to an elevated number of file operations within a time window characterized by malicious activity, in contrast to one with only benign activity. These observations are validated by Figures 2(b) and 2(f), which highlight the propensity of ransomware to execute a substantially higher number of operations on a file within a given time window. Lastly, the number of operations on the file complements the number of processes feature, introducing another valuable dimension for the classifier to consider.
@@ 5. Experimental Setup
5.1. Dataset and Ransomware Defenses
5.1.1. Dataset of Process Operations.
The File-Based Behavioral Detector module of Minerva is trained on features calculated from multiple different datasets. The data for benign and traditional ransomware were obtained from the authors of ShieldFS (Continella et al., 2016) and RWGuard (Mehnaz et al., 2018), enriched with traces from other traditional ransomware families. Regarding the evasive multiprocess ransomware, we utilize the evasive multiprocess ransomware families presented by De Gaspari et al. (De Gaspari et al., 2020) in all their evasive configurations. Moreover, the dataset on adaptive ransomware was obtained by altering the behavior of the ransomware introduced by the same authors, to manipulate the features used by Minerva for detection, as discussed in Section 5.2.2.
5.1.2. Extracted Features Dataset
From the above-mentioned dataset, we derive four separate feature dataset: benign feature dataset, traditional ransomware feature dataset, evasive multiprocess ransomware, and adaptive ransomware feature dataset. Minerva models are only trained on benign, traditional, and multiprocess feature dataset. We stress that the adaptive ransomware features dataset is never used in the training process. It is only used in the testing phase, to assess the resilience of Minerva to adaptive attacks and its ability to generalize to unseen ransomware.
5.1.3. Ransomware Defenses
In our evaluation, we compare our approach against two state-of-the-art ransomware detection approaches: ShieldFS (Continella et al., 2016) and RWGuard (Mehnaz et al., 2018). As highlighted in recent publications (De Gaspari et al., 2022a; Zhou et al., 2023), ShieldFS and RWGuard are still the reference benchmarks for behavioral-based ransomware detection.
5.2. Implementation Details
5.2.1. Behavioral Detector
Minerva uses an ensemble multi-tier architecture, where each tier’s classifier monitors file activity occurring within a specific time window. In our implementation, we use the following time windows for each tier: $[0.25,0.5,1,2,4]$ seconds. These time windows were selected to balance detection speed, system efficiency, and robustness. For each tier, we employ a fully connected neural network classifier.
5.2.2. Adaptive Ransomware
@@
To assess the resilience of Minerva against evasion attacks, we devise new families of adaptive ransomware specifically engineered to evade its detection. These adaptive ransomware are enhanced variants of the evasive multiprocess ransomware (De Gaspari et al., 2022a), improved to manipulate one or more of Minerva’s features to evade detection. We design three families of adaptive ransomware. (1) The adaptive R/W entropy ransomware adjusts the read and write entropy features by introducing artificial, low entropy read and write operations, thereby modulating the average entropy of operations within a given time window. We generate multiple variants of this adaptive family, with an up to $8$x reduction in average operation entropy. (2) The adaptive operation number ransomware alters the number of operations executed within a time window by increasing the size of each individual operation. We developed multiple variants of this family, resulting in up to $90\%$ reduction in number of operations. (3) The adaptive R/W ratio ransomware modifies the file read and write ratios within a specific window by slowing down the ransomware encryption activity. We create several variants for this family, with up to $80\%$ slowdown in the encryption speed.
We stress that all these adaptive families are not included in the training data, and are therefore unseen by the classifier.
@@ 6. Evaluation
The evaluation is structured as follows. Section 6.1 introduces the evaluation metrics. Section 6.2 analyzes Minerva’s detection time. Sections 6.36.46.5 present the detection performance of Minerva over various types of ransomware. Section 6.6 provides insights on the explainability of Minerva and how different features affect the classification outcome. Lastly, Section 6.7 discusses the system overhead introduced by Minerva’s modules.
@@ 6.1. Metrics
Benign Traditional Multiproc. Overall
Time window $TNR$ $TPR_{T}$ $TPR_{M}$ Accuracy Precision Recall F1
0.25sec 98.53 98.16 99.93 98.88 99.26 99.05 99.15
0.5sec 98.54 98.58 99.93 99.02 99.26 99.26 99.26
1sec 98.56 98.75 99.94 99.09 99.28 99.35 99.31
2sec 98.79 98.89 99.94 99.21 99.39 99.42 99.40
4sec 98.94 99.11 99.94 99.33 99.47 99.53 99.50
stacked 97.67 99.25 99.98 98.96 98.84 99.61 99.23
@@ 6.2. Detection Time
Detection time is a critical metric for assessing ransomware detection performance. From the ransomware perspective, hastening the malicious activity is paramount to achieving the objective of encrypting a maximum number of the victims’ files, thereby increasing the likelihood of ransom payment. Conversely, from the defender’s standpoint, minimizing the time required to detect the ongoing ransomware activity within the system reduces the probability of significant data loss. Furthermore, prompt detection of ransomware activity enables the adoption of near-zero overhead data recovery approaches, as we discuss in detail in Section 7.
@@
In Figure 3, we report TPR, TNR, and the time to detect the ransomware activity in the system by Minerva for varying numbers of $K$. As mentioned in Section 4, $K$ indicates the number of consecutive windows where malicious activity is detected by Minerva in order to classify the operations as malicious. Different values of $K$ provide a trade-off to balance between detection time and false positive rates. Figure 3 shows that increasing the value of $K$ leads to a corresponding increase in $TNR$, indicating better end-user experience due to the decreased occurrence of benign processes flagged as ransomware. This is expected behavior, as for a well-trained classifier, the likelihood of $K$ consecutive misclassifications is inversely proportional to the value of $K$. Conversely, the average detection time is directly proportional to $K$. Figure 3 shows the proportionality relation is linear, with the average detection time being almost exactly equal to the minimum window size times $K$. This is due to the high TPR achieved by the 0.25sec window classifier, as we discuss later in Section 6.3. Finally, the overall TPR is nearly unaffected by $K$, decreasing by ${\sim}0.4$ percentage points between $K=1$ and $K=5$. The motivation behind this behavior is twofold. First, the TPR of Minerva is consistently high across all windows with malicious activity, meaning that typically every window with ransomware activity is correctly flagged by Minerva. Therefore, given $K$ windows with malicious activity, the probability that at least one is flagged as benign is low. Second, as illustrated in Figure 4, the total duration of ransomware activity on a single file lasts less than 1 second for ${\sim}95\%$ of all files. This behavior effectively puts an upper bound on the number of maximum windows $k$ required for detection. In fact, if there are less than $K$ windows of activity on a given file, Minerva will still attempt classification using the available windows. For example, considering the $0.25sec$ window classifier, since ${\sim}95\%$ of all files have less than 1 second of ransomware activity, Minerva will use at most $4$ windows ($1sec/0.25sec$) to perform classification on these files, regardless of $K$. Therefore, increasing $K$ beyond a certain value impacts the classification only on a very small percentage of files.
@@
Based on this empirical evaluation, we found that $K=2$ provides the best trade-off between detection time, TNR, and TPR. For $K=2$, Minerva detects over 99% of the ransomware activity within $0.52sec$ ($0.25sec$ standard deviation) with low FPR, reducing the probability of significant data loss. Furthermore, as we discuss in Section 7, the fast detection rate of Minerva enables the use of near-zero overhead data recovery approaches, effectively nullifying data loss. Henceforth, unless explicitly stated otherwise, all subsequent experiments are presented for $K=2$.
@@ 6.3. Ransomware Detection Performance
This section evaluates the ransomware detection rate of Minerva and compares it to widely used benchmark papers in the area. We analyze the performance of our detector against two different categories of ransomware: traditional ransomware, and evasive multiprocess ransomware.
Approach Traditional $TPR$ Multiprocess $TPR$
ShieldFS (Continella et al., 2016) 100.00 0.00
RWGuard (Mehnaz et al., 2018) 98.45 0.00
Minerva (ours) 99.25 99.98
@@
Time window $TNR$ $TPR$ Accuracy Precision Recall F1

### v3

@@ Abstract.
This paper presents Minerva, a novel robust approach to ransomware detection. Minerva is engineered to be robust by design against evasion attacks, with architectural and feature selection choices informed by their resilience to adversarial manipulation. We conduct a comprehensive analysis of Minerva across a diverse spectrum of ransomware types, encompassing unseen ransomware as well as variants designed specifically to evade Minerva. Our evaluation showcases the ability of Minerva to accurately identify ransomware, generalize to unseen threats, and withstand evasion attacks. Furthermore, over $99\%$ of detected ransomware are identified within $0.52sec$ of activity, enabling the adoption of data loss prevention techniques with near-zero overhead.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 20th ACM ASIA Conference on Computer and Communications Security; August 25 – August 29, 2025; Hanoi, Vietnamisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Malware and its mitigationccs: Computing methodologies Machine learning
@@ 3. Threat Model
Trusted Components. We assume Minerva’s Disk Activity Monitor (DAM) and File-Based Behavioral Detector (FBD) components to be trusted, meaning that ransomware processes cannot tamper with their functionality. The DAM module is a driver running in kernel space, which is protected by the operating system. The FBD is a user-space component that communicates with the DAM through a protected, read-only communication port. The communication’s security context is set up so that only the system group associated with the FBD can access it, and the FBD runs with elevated (administrator) privileges to prevent tampering by other processes. We do not consider kernel-space ransomware, as malware with kernel-level access can effectively bypass any system-level countermeasure. These assumptions are in line with prior work in the field (Continella et al., 2016; Mehnaz et al., 2018; Kharraz and Kirda, 2017; Elkhail et al., 2023).
@@ 4.1. Minerva Architecture
From a high-level view, Minerva is comprised of two separate modules: (1) Disk Activity Monitor (DAM), and (2) File-Based Behavioral Detector (FBD). In what follows, we explain how these two modules work.
@@
The DAM module is a shim that provides system-wide, real-time information on all file system operations performed. We implemented the module as a driver that uses the I/O Request Packet (IRP) data structure provided by Windows systems to transparently log filesystem operations system-wide. Each filesystem-related system call performed by a process can result in many lower-level operations. The DAM captures and filters all these low-level operations before processing them into feature vectors, which are then forwarded to the FBD module of Minerva. To minimize overhead, we logically divide files into critical and non-critical: critical files encompass all user files that are high-priority targets for ransomware, while non-critical files comprise system and configuration files that can easily be recovered after an attack. Minerva automatically identifies as critical files all files located in non-system directories such as Documents, Downloads, and Desktop on Windows systems. System, program, temporary, and configuration files are excluded from monitoring. This includes all files in system directories (e.g., Windows) and in directories used by programs to store configuration and logging data (e.g., Program Files, ProgramData in Windows systems). Minerva’s critical files criteria can be further customized by system administrators to include or exclude additional directories and files. The DAM excludes any operation performed on non-critical files and computes per-file feature vectors for all other opened files.
@@
Figure 1 illustrates the architecture of the FBD module of Minerva. This module comprises an ensemble of machine-learning classifiers trained to identify ransomware activity by analyzing the behavioral profile of files. For each critical file in the system, Minerva computes several features based on the operations received by the file within a predefined time window. These features define that file’s behavioral profile for that time window, which the Detector uses to determine whether the file has undergone malicious activity within the window. To withstand variations in ransomware behavior, such as adaptive attacks aimed at slowing ransomware activity to span multiple windows, Minerva employs a multi-tier architecture. Each tier implements a distinct fully-connected Deep Neural Network classifier trained on feature vectors computed over progressively longer time windows: low-tier classifiers utilize feature vectors computed over short time windows (e.g., $0.25$ sec), while high-tier classifiers utilize feature vectors calculated over longer windows (e.g., $4$ sec). At the end of each time window, Minerva queries the classifier of the corresponding tier to detect malicious behavioral profiles. In order to confirm malicious activity on a file, Minerva requires that at least one of its classifiers reports $K$ consecutive windows with detected malicious activity. If a classifier reports malicious activity for $K$ consecutive windows, all processes that were active on the specific file during those windows are flagged as ransomware processes. This procedure is carried out for each active critical file in the system that is receiving read or write operations.
@@ 4.2. Feature Analysis
Minerva leverages seven distinctive file-level features that accurately describe the behavioral profile of files: read entropy, write entropy, write ratio, read ratio, number of processes, number of operations, and read/write ratio. We carefully selected these features following an analysis of a dataset containing both benign and ransomware processes to validate our initial intuition behind the choice. Our experimental evaluation demonstrates the significant challenge ransomware faces in modifying its behavior to successfully evade all of these features, given the contrastive nature between different features.
@@
The number of read and write operations that a file undergoes within a given time window offers essential context to weigh the significance of the other features. Benign processes commonly execute occasional file operations of relatively small size over their lifetime, leading to file-based behavioral profiles characterized by a low number of operations within each time window. This is apparent from Figures 2(b) and 2(f), which depict the distribution and CDF of the number of operations received by a file over 1-second intervals. Certain traditional ransomware favor using a limited number of large read and write operations during the encryption process, likely aiming to maximize I/O performance. Conversely, the majority of traditional ransomware utilize numerous small operations on the file, presumably to minimize the likelihood of detection. Similarly, in an effort to closely imitate the behavior of benign processes, evasive multiprocess ransomware imposes restrictions on the number of operations performed by each individual ransomware process and on the size of each operation (De Gaspari et al., 2022a; Zhou et al., 2023). The restriction imposed on individual operation size forces evasive multiprocess ransomware to use a greater overall number of read/write operations, distributing them across the various ransomware processes. In terms of file-based behavioral profiles, this behavior leads to an elevated number of file operations within a time window characterized by malicious activity, in contrast to one with only benign activity. These observations are validated by Figures 2(b) and 2(f), which highlight the propensity of ransomware to execute a substantially higher number of operations on a file within a given time window. Lastly, the number of operations on the file complements the number of processes feature, introducing another valuable dimension for the classifier to consider.
@@ 5. Experimental Setup
5.1. Dataset
5.1.1. Low-Level File Operations Data.
We utilize low-level file operation data (IRP logs) from three distinct types of ransomware: traditional, evasive multiprocess, and adaptive. Traditional ransomware includes operation data from 383 samples gathered in previous works (Continella et al., 2016; Mehnaz et al., 2018) — called Traditional Classic — along with data from 43 additional samples we collected from recent families (2023-2025) — called Traditional Current. Details on the collection process are provided in Appendix A.3. For evasive multiprocess ransomware, we incorporate operation data from (De Gaspari et al., 2020), covering all proposed evasive configurations. Adaptive ransomware operation data was generated by modifying the behavior of the evasive multiprocess ransomware to manipulate the features used by Minerva for detection, as detailed in Sections 5.2. Finally, for benign processes, we use operation data from (Continella et al., 2016), which includes extensive user activity recorded on 11 different machines across a diverse set of applications, including office suites, development tools such as Visual Studio, and archiving programs like WinRar. In total, the dataset comprises 476 ransomware samples and over 2000 benign applications. Full details are provided in Appendix A.2.
5.1.2. Train and Test Datasets
Minerva uses an ensemble multi-tier architecture, where each tier’s classifier monitors file activity occurring within a specific time window. Our implementation uses the following time windows for each tier: $[0.25,0.5,1,2,4]$ seconds. From the low-level file operations data, we compute Minerva features for each file and time window, deriving four separate datasets: (1) benign, (2) traditional ransomware, (3) evasive multiprocess ransomware, and (4) adaptive ransomware. We split the first three datasets in train and test based on a $70\%-30\%$ split and train Minerva on the three combined training datasets. The adaptive ransomware dataset is only used as a test set to assess the resilience of Minerva to adaptive attacks and not for training, with the exception of two R/W variants that are used to train Minerva on the relation between entropy and number of files operations, as detailed in Section 5.2. Unless explicitly stated otherwise, all reported results are based on this train-test split.
5.2. Adaptive Ransomware
@@
We devise three new families of adaptive ransomware specifically engineered to evade Minerva: (1) adaptive R/W entropy, (2) adaptive operation number, and (3) adaptive R/W ratio. These adaptive ransomware are enhanced versions of the mimicry evasive multiprocess ransomware (De Gaspari et al., 2022a), refined to adversarially manipulate one or more of Minerva’s detection features. (1) The adaptive R/W entropy ransomware adjusts the read and write entropy features by injecting artificial, low-entropy read and write operations, thereby modulating the average entropy of operations within a given time window. We generate six variants of this adaptive family, with entropy reductions ranging from $2$x up to $8$x. We train Minerva on the first two variants to help the classifiers learn the relationship between entropy and the number of operations. The remaining four variants, which remain unseen during training, are used to evaluate Minerva’s ability to generalize and detect previously unseen evasive strategies. (2) The adaptive operation number ransomware alters the number of operations executed within a time window by increasing the size of each individual operation. We developed nine variants of this family, reaching up to $90\%$ reduction in the number of operations. All variants are excluded from training. (3) The adaptive R/W ratio ransomware modifies the file read and write ratios within a specific window by slowing down the ransomware encryption activity. We generate six variants for this family, with up to $80\%$ slowdown in the encryption speed. All variants are unseen by the classifier during training.
@@ 6.1. Metrics
Benign Trad. Classic Trad. Current Multiproc. Overall
Window $TNR$ $TPR_{T}$ $TPR_{C}$ $TPR_{M}$ Acc. Prec. Recall F1
0.25sec 98.34 97.99 99.65 99.93 98.76 99.17 98.97 99.07
0.5sec 98.38 98.50 99.66 99.95 98.95 99.19 99.24 99.21
1sec 98.33 98.57 99.66 99.96 98.96 99.17 99.27 99.22
2sec 98.46 98.97 99.66 99.95 99.13 99.23 99.46 99.35
4sec 98.91 99.09 99.67 99.95 99.31 99.45 99.52 99.49
stacked 97.25 99.23 99.70 99.98 98.82 98.63 99.61 99.12
@@ 6.2. Detection Time
In Figure 3, we report TPR, TNR, and the time to detect the ransomware activity in the system by Minerva for varying numbers of $K$. As mentioned in Section 4, $K$ indicates the number of consecutive windows where malicious activity is detected needed for Minerva to classify the operations as malicious. Different values of $K$ provide a trade-off to balance between detection time and false positive rates. Figure 3 shows that increasing the value of $K$ leads to a corresponding increase in $TNR$, indicating better end-user experience due to the decreased occurrence of benign processes flagged as ransomware. This is expected, as the likelihood of $K$ consecutive misclassifications is inversely proportional to the value of $K$. Conversely, the average detection time is directly proportional to $K$. Figure 3 shows the proportionality relation is linear, with the average detection time being almost exactly equal to the minimum window size times $K$. This is due to the high TPR achieved by the 0.25sec window classifier, as we discuss later in Section 6.3. Finally, the overall TPR is nearly unaffected by $K$, decreasing by ${\sim}0.4$ percentage points between $K=1$ and $K=5$. The motivation behind this behavior is twofold. First, the TPR of Minerva is consistently high across all windows with malicious activity, meaning that typically every window with ransomware activity is correctly flagged by Minerva. Therefore, given $K$ windows with malicious activity, the probability that at least one is flagged as benign is low. Second, as illustrated in Figure 4, the total duration of ransomware activity on a single file lasts less than 1 second for ${\sim}95\%$ of all files. This behavior effectively puts an upper bound on the number of maximum windows $k$ used for detection. When there are $n<K$ consecutive windows of activity on a given file, Minerva will attempt classification using the $n$ available windows, regardless of $K$. Therefore, increasing $K$ beyond a certain value impacts the classification only on a very small percentage of files.
@@
Based on this empirical evaluation, we found that $K=2$ provides the best trade-off between detection time, TNR, and TPR. For $K=2$, Minerva detects over 99% of the ransomware activity within $0.52(\pm 0.25)sec$ with low FPR, reducing the probability of significant data loss. Furthermore, as we discuss in Section 7, the fast detection rate of Minerva enables the adoption of near-zero overhead data recovery approaches, effectively nullifying data loss. Henceforth, unless explicitly stated otherwise, all subsequent experiments are presented for $K=2$.
@@ 6.3. Ransomware Detection Performance
Window $TNR$ $TPR$ Acc. Prec. Recall F1